What is the equation of a line passing through the points ( 4,3 ) and ( 6,4 )
Answer by lwsshak3(11628)

What is the equation of a line passing through the points ( 4,3 ) and ( 6,4 )
***
Vertex form of equation for a line: y=mx+b, m=slope, b=y-intercept
For given line:
slope=∆y/∆x=(4-3)/6-4)=1/2
equation:
y=x/2+b
solve for b using coordinates(4,3) of given point on the line.
3=4/2+b
b=1
equation:
y=x/2+1